Assume Nortel Networks contracted to provide a customer with Internet infrastructure for $2,700,000. The project began in 2018 and was completed in 2019. Data relating to the contract are summarized below: 2019 2018 356,000 $2,135,000 Costs incurred during the year Estimated costs to complete as of 12/31 Billings during the year Cash collections during the year 1,424,000 475,000 1,760, 000 278,000 1,820,000 Required: 1. Compute the amount of revenue and gross profit or loss to be recognized in 2018 and 2019 assuming Nortel recognizes revenue over time according to percentage of completion.
Answer #1
2018 2019
Total cost to be incurred 356000+1424000=1780000 2135000
% of completion [total cost incurred till date /total cost to be incurred]

356000/1780000

.20 or 20%

2135000/2135000

100%

Gross profit 2700000-1780000=920000 2700000-2135000=565000
Revenue

[2700000*.20]-0

540000

[2700000*100%]-540000

2700000-540000

2160000

gross profit to be recognized

[920000*20%]- 0

184000-0

184000

[565000*100%]- 184000

565000-184000

381000

**Revenue recognized = [contract price * % of completion] - revenue previously recognized

**gross profit recognized = [ total gross profit *% of completion ] - gross profit previously recognized